EDITORIAL

Terry Goodman

Please take note of the subscriber survey found on page 157. We want to make sure MJMS reflects the preferences of the majority of our readers.

We want to use the editorial page in this issue to recognize our referees. We appreciate their conbribution to MJMS and look forward to working with them in the future.

List of Referees

Charles J. Allard, Polo R-VII High School Mohammad K. Azarian, University of Evansville John E. Bernard, University of Texas–Pan American Albert Dixon, College of the Ozarks John J. Edgell, Jr., Southwest Texas State University Vince Edmondson, Central Missouri State University Larry Eifler, University of Missouri–Kansas City Ekemezie J. Emeka, Columbia College Martin Erickson, Northeast Missouri State University Russell Euler, Northwest Missouri State University David Ewing, Central Missouri State University Joe Flowers, Northeast Missouri State University Jayanthi Ganapathy, University of Wisconsin–Oshkosh Ramesh Garimella, Northwest Missouri State University Victor Gummersheimer, Southeast Missouri State University K.L.D. Gunawardena, University of Wisconsin–Oshkosh Leon Hall, University of Missouri–Rolla Homer Hampton, Central Missouri State University Dennis R. Harmon, Missouri Southern State College Stan Hartzler, Northwest Missouri State University Troy Hicks, University of Missouri–Rolla Larry K. Johnson, Central Missouri State University N. J. Kuenzi, University of Wisconsin–Oshkosh Charles E. Lamb, University of Texas–Austin Yi–Hsin Liu, University of Nebraska–Omaha William Livingston, Missouri Southern State College Roger Marty, Cleveland State University Gary McDonald, Northwest Missouri State University

106

List of Referees (cont.)

Rhonda L. McKee, Central Missouri State University Kandasamy Muthvel, University of Wisconsin–Oshkosh Ollie Nanyes, Bradley University Krzysztof Ostaszewski, University of Louisville Bob Prielipp, University of Wisconsin–Oshkosh Don Redmond, Southern Illinois University Jau–Shyong Shiue, University of Nevada–Las Vagas Shing S. So, Central Missouri State University Songlin Tian, Central Missouri State University Mohan Tikoo, Southeast Missouri State University Alvin Tinsley, Central Missouri State University Bobby Neal Winters, Pittsburg State University